OpenStudy (anonymous):

All of the following expressions are equivalent except _____. a) 2 -x b) x- 2 c) -2 +x d) x+ (-2)

OpenStudy (anonymous):

2-x

OpenStudy (anonymous):

b=x-2 arrange c=x-2 d=x-2

OpenStudy (anonymous):